# Searching for Planets Orbiting Vega with the James Webb Space Telescope

### v1

@@ 1 Introduction
Section 2 describes the observations and 3 the data reduction steps. 4 presents images close to the star ( 10″) and over a much wider field, describes the limiting contrast level (3.4), lists a number of sources detected in the images, and details a search for scattered disk emission (3.7). 5 discusses both the sub-Jovian mass limits achieved here in the context of planet formation mechanisms and the limits on scattered light.
@@ 2 Observations
JWST observed Vega through the round M430R coronagraphic mask simultaneously at two wavelengths (F210M and F444W), one each in the Long and Short wavelengths arms of NIRCam (Rieke et al. 2023; Table 2). The observations were obtained at two roll angles separated by 7°. Because of the brightness of Vega, we split the observations into: 1) a set of subarray (SUB320) exposures optimized to avoid saturation close to the star but with a limited field of view extending only 10″; and 2) a set of observations in full array mode providing coverage to the full extent of the debris disk out beyond 30″ (Su et al., 2013; Matrà et al., 2020).
@@ 3.5 Point Sources Interior or Close to the Debris Disk
Outside the speckle-dominated region (Figure 2) we identified two sources (S1 and S2) approximately 6″ (46 au) away from Vega (Table 3, Figure A1). Both appear to be extended. The first source is also identified with an object seen in MIRI data. A third, marginal 2.5-3 detection is described in the Appendix.
@@ 3.7 Limits to Disk Emission
Although Vega has a prominent debris disk seen in thermal emission, detection of a visible light counterpart has proven elusive particularly as the disk is nearly face-on. A deep HST program to search for scattered light around Vega (PID 16666) shows tentative evidence for scattered light signal from 10.″5 - 30″ with a peak surface brightness MJy sr-1 and a relatively shallow radial profile (Wolff et al, 2024). Detection of a disk in the near-IR is more challenging than at visible wavelengths due to the lower stellar flux () and typically smaller scattering cross-sections for small grains (). To search for scattered light we used a PSF-subtracted image (Figure 5a), computed via a simple RDI, similar to what is described in Sec. 3.3 but without using the other roll to avoid self-subtraction of the disk. The 5 available reference frames (from the 5-POINT dither pattern described in Sec. 2) were combined in a linear combination, the coefficients of which were computed by maximizing the speckle subtraction in the inner 100100 pixel around the center of the star.
@@ 4.3 Limits to Scattered Disk Light
The detection of scattered light around Vega is challenging due to a number of factors. First, the disk will be close to face-on which minimizes the integrated optical depth. Second, at NIRCam wavelengths, the stellar brightness is fainter and the scattering cross sections typically lower than in the visible. To compare the NIRCam results with the visible detection (or limit; Wolff et al (2024)), we approximate the scattered light surface brightness, at a distance from Vega as:
where , is the scattering optical depth, is the distance to the star from Earth, and is the measured flux density of the star at F444W. For the 0.5 MJy sr-1 limits estimated here (Figure 5), the limits to the optical depth at 10″-20″ are which can be compared to the tentative HST detection (Wolff et al 2024, submitted) which suggests at visible wavelengths. The MIRI-determined emission optical depths are at comparable separations (Su et al., 2024). For small “astronomical” silicate grains, 0.4 (0.16) m, the ratio of absorption to scattering cross-sections is (Draine & Lee, 1984), so that the emission optical depth determined from the MIRI data would imply a scattering optical depth of at F444W. Smaller grains would lead to considerably smaller predicted scattering optical depths, relative to the upper limits measured here of a few . Similar arguments apply even if the grains are composed of ices rather than by astronomical silicates as has been suggested to explain HST scattered light observations (Wolff et al, 2024; Tazaki et al., 2021). Thus, on the basis of the HST and MIRI data and typical grain properties, the failure to detect scattered light at F444W is not surprising.
@@ 5.2.1 Summary of Contrast and Mass Limits
If we extend this distribution down to 0.5 for the full range of separations (10-100 AU; black curve in Figure 10b), then the expected frequency increases to unity probability of finding a planet at this mass sensitivity level. However, NIRCam achieves its best mass sensitivity only outside of 10″ (50 au depending on the distance to the target; Figure 9) and the expected increase in the planet occurrence rate is about a factor of 4, from 8.9% to 30% (lower, blue curve in Figure 10).
@@ 6 Conclusions
This source of marginal significance, denoted , is located 1.4” (10.7 au) South of Vega Figure A2. It has no counterpart at F210M or in the MIRI data. If were associated with Vega, then its apparent magnitude of [F444W]=16.5 mag would correspond to an absolute mag of 17.1 mag and a mass of approximately 2-3MJup based on a Sonora Bobcat model (Marley et al., 2021) with an age of 700 Myr and effective temperature of 200-250 K. Such an object would not be detectable in our F210M data.

### v2

@@ 1 Introduction
Section 2 describes the observations. 3 describes the data reduction steps, presents images close to the star ( 10″) and over a much wider field, describes the limiting contrast level (3.4), lists a number of sources detected in the images, and details a search for scattered disk emission (3.7). 4 assesses the potential of the point sources as exoplanet candidates and discusses the surface brightness limits in terms of disk properties. 5 discusses both the sub-Jovian mass limits achieved here in the context of planet formation mechanisms and the limits on scattered light. 6 gives our conclusions.
@@ 2 Observations
JWST observed Vega through the round M430R coronagraphic mask simultaneously at two wavelengths (F210M and F444W), one each in the Long and Short wavelengths arms of NIRCam (Rieke et al. 2023; Girard et al. 2022; Table 2). The observations were obtained at two roll angles separated by 7°. Because of the brightness of Vega, we split the observations into: 1) a set of subarray (SUB320) exposures optimized to avoid saturation close to the star but with a limited field of view extending only 10″; and 2) a set of observations in full array mode providing coverage of the full extent of the debris disk out beyond 30″ (Su et al., 2013; Matrà et al., 2020).
@@ 3.5 Point Sources Interior or Close to the Debris Disk
Outside the speckle-dominated region (Figure 4) we identified two sources (S1 and S2) approximately 6″ (46 au) away from Vega (Table 3, Figure A1). Both appear to be extended. The first source is also identified with an object seen in MIRI data (4.1). A third, marginal 2.5-3 detection is described in the Appendix (Figure A2, Table A1).
@@ 3.7 Limits to Disk Emission
Although Vega has a prominent debris disk seen in thermal emission, detection of a visible light counterpart has proven elusive particularly as the disk is nearly face-on. A deep HST program to search for scattered light around Vega (PID 16666) shows tentative evidence for scattered light signal from 10.″5 - 30″ with a peak surface brightness MJy sr-1 and a relatively shallow radial profile (Wolff et al., 2024). Detection of a disk in the near-IR is more challenging than at visible wavelengths due to the lower stellar flux () and typically smaller scattering cross-sections for small grains (). To search for scattered light we used a PSF-subtracted image (Figure 5a), computed via a simple RDI, similar to what is described in Sec. 3.3 but without using the other roll to avoid self-subtraction of the disk. The 5 available reference frames (from the 5-POINT dither pattern described in Sec. 2) were combined in a linear combination, the coefficients of which were computed by maximizing the speckle subtraction in the inner 100100 pixel around the center of the star.
@@ 4.3 Limits to Scattered Disk Light
The detection of scattered light around Vega is challenging due to a number of factors. First, the disk will be close to face-on which minimizes the integrated optical depth. Second, at NIRCam wavelengths, the stellar brightness is fainter and the scattering cross sections typically lower than in the visible. To compare the NIRCam results with the visible detection (or limit; Wolff et al. (2024)), we approximate the scattered light surface brightness, at a distance from Vega as:
where , is the scattering optical depth, is the distance to the star from Earth, and is the measured flux density of the star at F444W. For the 0.5 MJy sr-1 limits estimated here (Figure 5), the limits to the optical depth at 10″-20″ are which can be compared to the tentative HST detection (Wolff et al 2024, submitted) which suggests at visible wavelengths. The MIRI-determined emission optical depths are at comparable separations (Su et al., 2024). For small “astronomical” silicate grains, 0.4 (0.16) m, the ratio of absorption to scattering cross-sections is (Draine & Lee, 1984), so that the emission optical depth determined from the MIRI data would imply a scattering optical depth of at F444W. Smaller grains would lead to considerably smaller predicted scattering optical depths, relative to the upper limits measured here of a few . Similar arguments apply even if the grains are composed of ices rather than by astronomical silicates as has been suggested to explain HST scattered light observations (Wolff et al., 2024; Tazaki et al., 2021). Thus, on the basis of the HST and MIRI data and typical grain properties, the failure to detect scattered light at F444W is not surprising.
@@ 5.2.1 Summary of Contrast and Mass Limits
If we extend this distribution down to 0.5 for the full range of separations (10-100 AU; black curve in Figure 10), then the expected frequency increases to unity probability of finding a planet at this mass sensitivity level. However, NIRCam achieves its best mass sensitivity only outside of 10″ (50 au depending on the distance to the target; Figure 9) and the expected increase in the planet occurrence rate is about a factor of 4, from 8.9% to 30% (lower, blue curve in Figure 10).
@@ 6 Conclusions
This source of marginal significance, denoted , is located 1.4” (10.7 au) South of Vega Figure A2. It has no counterpart at F210M or in the MIRI data. If were associated with Vega, then its apparent magnitude of [F444W]=16.5 mag (Table A1) would correspond to an absolute mag of 17.1 mag and a mass of approximately 2-3MJup based on a Sonora Bobcat model (Marley et al., 2021) with an age of 700 Myr and effective temperature of 200-250 K. Such an object would not be detectable in our F210M data.